# vesselFM: A Foundation Model for Universal 3D Blood Vessel Segmentation

### v1

@@ 4 Experiments and Results
In this section, we elaborate on our findings and showcase vesselFM’s performance on three tasks: zero-shot, one-shot, and few-shot segmentation. We evaluate vesselFM on four 3D blood vessel segmentation datasets of unseen clinically (SMILE-UHURA [9], MSD8 [2]) and pre-clinically (OCTA [54, 17], BvEM [51]) relevant domains. In this context, we extract three patches of shape $\text{128}^{\text{3}}$ from each of these evaluation datasets and use the remaining data for testing and validation (see Suppl. 6 for more details). With the three extracted patches, we define the one- and few-shot segmentation task as fine-tuning models on either one or all three patches. For zero-shot evaluation, we apply models out-of-the-box on the test data without prior fine-tuning. This setup mimics clinical scenarios where annotated data is typically scarce.
We compare vesselFM with four foundation models designed for 3D medical image segmentation, including the generalizable 3D blood vessel segmentation model tUbeNet [24], the CT-specific VISTA3D [23], and the two SAM-like general-purpose segmentation models SAM-Med3D [53] and MedSAM-2 [61]. In our experiments, we exclude the classes of the four datasets used for evaluation (see Table 1, upper section) from the curation of $\mathcal{D}_{\text{real}}$ and $\mathcal{D}_{\text{flow}}$. This results in $\mathcal{D}_{\text{real}}$ containing 90,685 labeled patches. We generate $\mathcal{D}_{\text{flow}}$ by sampling 10,000 image-mask pairs from $\mathcal{F}$ on a single RTX A6000 GPU over the course of three days. To curate $\mathcal{D}_{\text{drand}}$, we sample 500,000 image-mask pairs from our domain randomized generative pipeline. All images-mask pairs are of shape $\text{128}^{\text{3}}$. We train vesselFM using all three data sources, with weights assigned roughly according to their sizes (70% $\mathcal{D}_{\text{drand}}$, 20% $\mathcal{D}_{\text{real}}$, and 10% $\mathcal{D}_{\text{flow}}$). We opt for the UNet architecture proposed by Isensee et al. [26] to present our segmentation model. For flow matching, we use the UNet from Med-DDPM [14] to represent the learned velocity field $v$. We set the total number of time steps $N$ in (3) to 100. Following common practices, we report Dice scores and topology-aware centerline Dice (clDice) scores [43], which assess the preservation of tubular appearance and connectivity of blood vessels. Further details on our experimental setup can be found in Suppl. 9.
@@ 4.1 Quantitative and Qualitative Results
VesselFM exhibits exceptional zero-shot generalization on all four datasets, which cover a diverse array of unseen domains and even imaging modalities (e.g., OCTA and vEM). Surprisingly, vesselFM scores 5.86 Dice points higher than VISTA3D on MSD8, even though VISTA3D was trained on 11,454 CT volumes, including data from MSD8 itself. This highlights vesselFM’s strong inductive bias, enabled by training on our three proposed heterogeneous blood vessel data sources. Moreover, vesselFM outperforms the generalizable 3D blood vessel segmentation model tUbeNet, trained on four blood vessel datasets of varying imaging modalities. We observe that tUbeNet struggles in more complex imaging modalities where blood vessels do not have a stark contrast to background tissues (e.g., BvEM and MSD8). The general-purpose segmentation models SAM-Med3D and MedSAM-2 both fail to segment blood vessels in the zero-shot setting. Notably, vesselFM’s zero-shot results surpass few-shot results achieved by baseline models on SMILE-UHURA in Dice and clDice.
@@
Fine-tuning vesselFM in a one- or few-shot manner generally increases segmentation performance. Given that baseline models rely on architectures, which easily overfit to the small amount of training data provided in the one- and few-shot settings, we additionally compare vesselFM to a variant of the same configuration that is trained from scratch without being pre-trained on our three proposed data sources (see footnote in Table 2). We observe that neglecting our three proposed data sources causes a notable decrease in Dice and clDice scores, validating our rationale.
@@ 4.2 Ablation Studies
First, we ablate the relevance of our three proposed data sources, $\mathcal{D}_{\text{real}}$, $\mathcal{D}_{\text{drand}}$, and $\mathcal{D}_{\text{flow}}$ (see Table 3). To this end, we train vesselFM on each data source individually and progressively augment $\mathcal{D}_{\text{real}}$ with $\mathcal{D}_{\text{drand}}$ and $\mathcal{D}_{\text{flow}}$. A more complete ablation of our data sources covering all four evaluation datasets and all tasks can be found in Suppl. 8.
@@
Third, we ablate vesselFM’s segmentation model by experimenting with relevant medical 3D segmentation architectures (see Table 5).
@@
Our employed UNet variant [26] surpasses transformer-based [20, 22, 21] and ConvNeXt [34]-based [30, 42] architectures alike, which accurately represents the current landscape in medical image segmentation.
@@ 5 Conclusion and Outlook
In this work, we propose vesselFM, a foundational model for universal 3D blood vessel segmentation. VesselFM is capable of accurately segmenting 3D vasculature in previously unseen modalities and tissue types in zero-, one-, and few-shot settings, outperforming state-of-the-art medical image segmentation foundational models across all datasets and tasks by a wide margin. We enable zero-shot generalization by training vesselFM on three proposed heterogeneous data sources ($\mathcal{D}_{\text{real}}$, $\mathcal{D}_{\text{drand}}$, and $\mathcal{D}_{\text{flow}}$), which we extensively ablate in our experiments. Given that vesselFM pushes the frontier in the (pre-)clinically relevant task of 3D blood vessel segmentation, we hope that our work will enable more precise analysis of vascular disorders and foster more accurate diagnostic tools, ultimately resulting in improved patient outcomes. We advise future work to broaden the scope of vesselFM to other tubular structures (e.g., neurons, airways, or spinal cords) and experiment with tailored post-processing steps to improve blood vessel connectivity, potentially at the graph level [55]. Given that our foundation model is currently constrained to binary blood vessel segmentation, which may be viewed as a limitation, we further aim to extend it to multi-class segmentation tasks.
@@ Acknowledgments:
6 Pre-Processing of Datasets in $\mathcal{D}_{\text{real}}$
@@
We aim to ensure that all 23 datasets used to curate $\mathcal{D}_{\text{real}}$ (see Table 1) adhere to general vascular imaging characteristics and comply with our label quality standards. To this end, we apply carefully selected pre-processing steps (see Table 1, last column). First, we resample the TubeTK [7], SMILE-UHURA [9], TopCoW [59], MSD8 [2], tUbeNet [24], and DeepVesselNet [46] datasets, establishing appropriate blood vessel scales (e.g., a single vessel should not occupy 90% of the patch) and near-isotropic voxel sizes. In the case of the BvEM [51] dataset, only the labels are resampled, as the annotations were likely made on a downsampled version of the volume. As resampling may introduce label artifacts, we subsequently smooth affected labels using a median filter with a kernel size of $\text{5}^{\text{3}}$ followed by thresholding (threshold at 0.5). Since the 3D-IRCADb-01 [44] dataset contains labels of multiple structures, we solely keep venous system, artery, and portal vein labels, converting them to binary labels. The original labels of the HR-Kindney [29] dataset are of relatively poor quality. However, enabled by the high signal-to-noise ratio of the volume representing the HR-Kindney dataset, we improve label quality by applying Algorithm 1. To further enhance the visibility of blood vessels, the intensities of the VesSAP [48] and LS [5] datasets are clipped at the 2% and 98% percentiles, while the intensities of the MSD8 dataset are clipped at the 20% and 98% percentiles. We crop the MSD8 and 3D-IRCADb-01 datasets to retain only foreground structures, as the images otherwise would primarily consist of non-annotated anatomical structures. Additionally, we crop the borders from the BvEM volume, given that they predominantly contain artifacts.
@@
$\text{Mask}\leftarrow\text{Mask}\bullet\mathbf{1}_{5\times 5\times 5}$triangleright Close small gaps.
@@
From each evaluation dataset (see Table 1, upper section), we extract three patches of size $\text{128}^{\text{3}}$ for fine-tuning models in the one- and few-shot settings and use the remaining data for testing and validation. OCTA [54, 17]: We allocate three of the six samples provided in the OCTA dataset for one- and few-shot fine-tuning, and reserve the remaining three for model evaluation: two samples for testing and one sample for validation. The three samples used for fine-tuning are center-cropped to adhere to our target shape of $\text{128}^{\text{3}}$. SMILE-UHURA [9]: The fourteen samples in the SMILE-UHURA dataset are divided into one for validation, ten for testing, and three for extracting patches for fine-tuning. For the extraction of the fine-tuning patches, we pay special attention to extract patches highly representative of the characteristics of MRA scans contained in the SMILE-UHURA dataset (e.g., they contain vasculature, brain tissue, skull, and gyri/sulci). MSD8 [2]: We split the MSD8 dataset into one validation, 296 test, and three samples utilized to extract patches for fine-tuning. The patches for fine-tuning are chosen from representative regions and padded, if necessary, minimally in the $z$-dimension using reflective padding to conform to the target shape of $\text{128}^{\text{3}}$. BvEM [51]: The BvEM dataset contains solely a single volume of shape $\text{3571}\times\text{5145}\times\text{2496}$. We choose the first 130 slices to extract three $\text{128}^{\text{3}}$ patches for fine-tuning and one for validation. Then, we leave a buffer of 120 slices to minimize information leakage between the patches used for fine-tuning and testing. Lastly, we extract three bigger test volumes of shape $\text{500}^{\text{3}}$ from the remaining volume, limiting the overlap of these volumes with the fine-tuning patches in the $x$- and $y$-position as much as possible.
7 Parametrization for Domain Randomization
@@
8 More Detailed Ablation of Data Sources
@@
9 More Details on Experimental Setup
Training of vesselFM.
We employ a combination of Dice and cross-entropy loss functions, weighted by 0.9 and 0.1, respectively. VesselFM is trained on a single V100 GPU (32GB) with a batch size of 8 until convergence. The learning rate is set to $\text{10}^{\text{-4}}$. We utilize linear warm-up and a cosine annealing learning rate decay to $\text{10}^{\text{-6}}$. During training, we sample classes near uniformly from $\mathcal{D}_{\text{real}}$ and also from $\mathcal{D}_{\text{flow}}$. We solely apply data augmentation to samples from $\mathcal{D}_{\text{real}}$. To optimize training efficiency, we perform data augmentation offline. Specifically, we apply, after extracting $\text{128}^{\text{3}}$ patches, random flipping and rotation (angle in [0, 45]) along all axes. Further, we perform random zooming (factor in [0.9, 1.3]) and random elastic deformation ($\sigma$ in [10, 20] and magnitude in [100, 500]). Finally, we scale intensities to the range of [0, 1].
Figure 11: More slices of exemplary images $\mathcal{X}_{\text{drand}}$ (a) and $\mathcal{X}_{\text{flow}}$ (b) with matching masks $\mathcal{M}_{\text{syn}}$ shown in translucent red.
One- and few-shot fine-tuning.
In the one- and few-shot setting, we fine-tune vesselFM using a similar setup with an initial learning rate of $\text{10}^{\text{-5}}$ decayed to $\text{10}^{\text{-7}}$. We train vesselFM until convergence, selecting the checkpoint with the best Dice scores on the respective validation volume. We apply lightweight data augmentations on the fly: random zooming (factor in [1, 1.3]), random shearing (shearing factors in [0, 0.4]), random flipping along all axes, random Gaussian noise ($\mu$ of 0.3, $\sigma$ in [0, 0.5]), random Gaussian smoothing ($\sigma$ in [0, 0.5]), and random histogram shifting (number of points in [5, 10]), followed by intensity scaling.
10 Additional Qualitative Zero-Shot Results

### v2

@@ 4 Experiments and Results
In this section, we elaborate on our findings and showcase vesselFM’s performance on three tasks: zero-shot, one-shot, and few-shot segmentation. We evaluate vesselFM on four 3D blood vessel segmentation datasets of unseen clinically (SMILE-UHURA [9], MSD8 [2]) and pre-clinically (OCTA [54, 17], BvEM [51]) relevant domains. In this context, we extract three patches of shape $\text{128}^{\text{3}}$ from each of these evaluation datasets and use the remaining data for testing and validation (see Suppl. A for details). With the three extracted patches, we define the one- and few-shot segmentation task as fine-tuning models on either one or all three patches. For zero-shot evaluation, we apply models out-of-the-box on the test data without prior fine-tuning. This setup mimics clinical scenarios where annotated data is scarce.
We compare vesselFM with four foundation models designed for 3D medical image segmentation: the generalizable 3D blood vessel segmentation model tUbeNet [24], the CT-specific VISTA3D [23], and the two SAM-like general-purpose segmentation models SAM-Med3D [53] and MedSAM-2 [61]. In our experiments, we exclude the classes of the four datasets used for evaluation (see Table 1, upper section) from the curation of $\mathcal{D}_{\text{real}}$ and $\mathcal{D}_{\text{flow}}$. We generate $\mathcal{D}_{\text{flow}}$ by sampling 10,000 image-mask pairs from $\mathcal{F}$ on a single RTX A6000 GPU over the course of three days. To curate $\mathcal{D}_{\text{drand}}$, we sample 500,000 image-mask pairs from our domain randomized generative pipeline. All images-mask pairs are of shape $\text{128}^{\text{3}}$. We train vesselFM using all three data sources, with weights assigned roughly according to their sizes (70% $\mathcal{D}_{\text{drand}}$, 20% $\mathcal{D}_{\text{real}}$, and 10% $\mathcal{D}_{\text{flow}}$). We opt for MONAI’s reimplementation of the UNet architecture proposed by Isensee et al. [26] to present our segmentation model. For flow matching, we use the UNet from Med-DDPM [14] to represent the learned velocity field $v$. We set the total number of time steps $N$ in (3) to 100. Following common practices, we report Dice scores and topology-aware centerline Dice (clDice) scores [43], which assess the preservation of tubular appearance and connectivity of blood vessels. Further details on our experimental setup can be found in Suppl. C.
@@ 4.1 Quantitative and Qualitative Results
VesselFM exhibits exceptional zero-shot generalization on all four datasets, which cover a diverse array of unseen domains and even imaging modalities (OCTA and vEM). Surprisingly, vesselFM scores 5.86 Dice points higher than VISTA3D on MSD8, even though VISTA3D was trained on 11,454 CT volumes, including data from MSD8 itself. This highlights vesselFM’s strong inductive bias, enabled by training on our three proposed heterogeneous blood vessel data sources. Moreover, vesselFM outperforms the generalizable 3D blood vessel segmentation model tUbeNet, trained on four blood vessel datasets of varying imaging modalities. We observe that tUbeNet struggles in more complex imaging modalities where blood vessels do not have a stark contrast to background tissues (e.g., BvEM and MSD8). The general-purpose segmentation models SAM-Med3D and MedSAM-2 both fail to segment blood vessels in the zero-shot setting. Notably, vesselFM’s zero-shot results surpass few-shot results achieved by baseline models on SMILE-UHURA in Dice and clDice.
@@
Fine-tuning vesselFM in a one- or few-shot manner generally increases segmentation performance. Given that some baseline models rely on networks, which may easily overfit to the small amount of training data provided in the one- and few-shot settings, we additionally compare vesselFM to a variant of the same configuration that is trained from scratch without being pre-trained on our three proposed data sources (see footnote in Table 2; see Suppl. D for more insights). We observe that neglecting our three proposed data sources causes a notable decrease in Dice and clDice scores, validating our rationale.
@@ 4.2 Ablation Studies
First, we ablate the relevance of our three proposed data sources, $\mathcal{D}_{\text{real}}$, $\mathcal{D}_{\text{drand}}$, and $\mathcal{D}_{\text{flow}}$ (see Table 3). To this end, we train vesselFM on each data source individually and progressively augment $\mathcal{D}_{\text{real}}$ with $\mathcal{D}_{\text{drand}}$ and $\mathcal{D}_{\text{flow}}$. A more complete ablation of our data sources covering all four evaluation datasets and all tasks can be found in Suppl. E.
@@
Third, we ablate vesselFM’s segmentation model by experimenting with relevant medical 3D segmentation networks (see Table 5).
@@
Our employed UNet variant surpasses transformer-based [20, 22, 21] and ConvNeXt [34]-based [30, 42] networks alike, which accurately represents the current landscape in medical image segmentation.
@@ 5 Conclusion and Outlook
In this work, we propose vesselFM, a foundational model for universal 3D blood vessel segmentation. VesselFM is capable of accurately segmenting 3D vasculature in previously unseen modalities and tissue types and performs superior to state-of-the-art medical image segmentation foundational models. We enable zero-shot generalization by training vesselFM on three proposed heterogeneous data sources ($\mathcal{D}_{\text{real}}$, $\mathcal{D}_{\text{drand}}$, and $\mathcal{D}_{\text{flow}}$), which we extensively ablate in our experiments. Given that vesselFM pushes the frontier in the (pre-)clinically relevant task of 3D blood vessel segmentation, we hope that our work enables novel insights into vascular disorders and fosters the development of advanced diagnostic tools, ultimately resulting in improved patient outcomes (see Suppl. F for discussion on clinical utility). We advise future work to experiment with tailored post-processing steps to improve blood vessel connectivity (potentially at the graph level [55]) and extend vesselFM to multi-class or instance segmentation tasks.
@@ Acknowledgments:
Appendix A Pre-Processing of Datasets in $\mathcal{D}_{\text{real}}$
@@
We aim to ensure that all 23 datasets used to curate $\mathcal{D}_{\text{real}}$ (see Table 1) adhere to general vascular imaging characteristics and comply with our label quality standards. To this end, we apply carefully selected pre-processing steps (see Table 1, last column). First, we resample the BvEM [51], TubeTK [7], tUbeNet [24], TopCoW [59], and DeepVesselNet [46] datasets, establishing appropriate blood vessel scales (e.g., a single vessel should not occupy 90% of the patch) and near-isotropic voxel sizes. In the case of the BvEM dataset, only the labels are resampled, as the annotations were likely made on a downsampled version of the volume. As resampling may introduce label artifacts, we subsequently smooth affected labels (TubeTK and tUbeNet) using Gaussian smoothing followed by thresholding. Since the 3D-IRCADb-01 [44] dataset contains labels of multiple structures, we solely keep venous system, artery, and portal vein labels, converting them to binary labels. The original labels of the HR-Kindney [29] dataset are of relatively poor quality. However, enabled by the high signal-to-noise ratio of the volume representing the HR-Kindney dataset, we improve label quality by applying Algorithm 1. To further enhance the visibility of blood vessels, the intensities of the VesSAP [48] and LS [5] datasets are clipped at the 0% and 98% percentiles, while the intensities of the MSD8 dataset are clipped at the 20% and 98% percentiles. We crop the MSD8 and 3D-IRCADb-01 datasets to retain only foreground structures, as the images otherwise would primarily consist of non-annotated anatomical structures. Additionally, we crop the borders from the BvEM volume, given that they predominantly contain artifacts.
@@
$\text{Mask}\leftarrow\text{Mask}\bullet\mathbf{1}_{3\times 3\times 3}$triangleright Close small gaps.
@@
From each evaluation dataset (see Table 1, upper section), we extract three patches of size $\text{128}^{\text{3}}$ for fine-tuning models in the one- and few-shot settings and use the remaining data for testing and validation. OCTA [54, 17]: We allocate three of the six samples provided in the OCTA dataset for one- and few-shot fine-tuning and reserve the remaining three for model evaluation: two samples for testing and one sample for validation. The three samples used for fine-tuning are center-cropped to adhere to our target shape of $\text{128}^{\text{3}}$. SMILE-UHURA [9]: The fourteen samples in the SMILE-UHURA dataset are divided into one for validation, ten for testing, and three for extracting patches for fine-tuning. For the extraction of the fine-tuning patches, we pay special attention to extract patches highly representative of the characteristics of MRA scans contained in the SMILE-UHURA dataset (e.g., they contain vasculature, brain tissue, skull, and gyri/sulci). MSD8 [2]: We split the MSD8 dataset into one validation sample, 296 test samples, and three samples utilized to extract patches for fine-tuning. The patches for fine-tuning are chosen from representative regions and padded, if necessary, minimally in the $z$-dimension using reflective padding to conform to the target shape of $\text{128}^{\text{3}}$. BvEM [51]: The BvEM dataset contains solely a single volume of shape $\text{3571}\times\text{5145}\times\text{2495}$. We choose the first 130 slices to extract three $\text{128}^{\text{3}}$ patches for fine-tuning and one for validation. Then, we leave a buffer of 120 slices to minimize information leakage between the patches used for fine-tuning and testing. Lastly, we extract three bigger test volumes of shape $\text{500}^{\text{3}}$ from the remaining volume, limiting the overlap of these volumes with the fine-tuning patches in the $x$- and $y$-position as much as possible.
Appendix B Parameters of Domain Randomization
@@
Appendix C More Details on Experimental Setup
Parametrization of segmentation model.
As already stated in the main manuscript, we opt for MONAI’s re-implementation of [26]’s UNet architecture, called DynUNet, to present our segmentation model. To be specific, we set strides to [[1, 1, 1], [2, 2, 2], [2, 2, 2], [2, 2, 2], [2, 2, 2], [2, 2, 2]], kernel_size to [[3, 3, 3], [3, 3, 3], [3, 3, 3], [3, 3, 3], [3, 3, 3], [3, 3, 3]], upsample_kernel_size to [[2, 2, 2], [2, 2, 2], [2, 2, 2], [2, 2, 2], [2, 2, 2]], filters to [32, 64, 128, 256, 320, 320], and activated residual connection-based convolution block (res_block). We found that this parametrization performed best in our experiments.
Training of vesselFM.
We employ a combination of Dice and cross-entropy loss functions, weighted by 0.9 and 0.1. VesselFM is trained on a single V100 GPU (32GB) with a batch size of 8 until convergence. The learning rate is set to $\text{10}^{\text{-4}}$. We utilize linear warm-up and a learning rate decay to $\text{10}^{\text{-5}}$. During training, we sample classes near uniformly from $\mathcal{D}_{\text{real}}$ and also from $\mathcal{D}_{\text{flow}}$. We solely apply data augmentation to samples from $\mathcal{D}_{\text{real}}$. To optimize training efficiency, we perform data augmentation offline. Specifically, we apply, after extracting $\text{128}^{\text{3}}$ patches, random flipping and rotation (angle in [0, 10]) along all axes followed by random elastic deformation ($\sigma$ in [10, 20] and magnitude in [100, 500]) and random zooming (factor in [0.9, 1.3]).
One- and few-shot fine-tuning.
In the one- and few-shot setting, we fine-tune vesselFM using a similar setup with a learning rate of $\text{10}^{\text{-5}}$. We train vesselFM until convergence, selecting the checkpoint with the best Dice score on the respective validation volume. We apply lightweight data augmentations on the fly: random zooming (factor in [1, 1.3]), random shearing (shearing factors in [0, 0.4]), random flipping, random Gaussian smoothing ($\sigma$ in [0, 0.5]), random Gaussian noise ($\mu$ of 0.3, $\sigma$ in [0, 0.01]), and random histogram shifting (number of points in [5, 10]).
Appendix D VesselFM From Scratch
’VesselFM (from scratch)’ (see Table 2) demonstrates a relatively strong performance compared to other baselines and even outperforms them on the OCTA and BvEM datasets in some metrics. We attribute this to their unique imaging artifacts and intensity patterns, which, unlike those in CT and MRA, have not been observed by any baseline during pre-training. Note that the human retinal OCTA sample present in tUbeNet’s training dataset exhibits poor label quality and differs significantly from the characteristics [54] of the murine cerebral OCTA images used in our experiments. Further, architectural biases (SAM-Med3D and MedSAM-2 rely on large Transformers) and mismatches in already learned representations (VISTA3D is exclusively trained on CT images; tUbeNet is exclusively trained on images with stark contrast to background tissues) may impede few-shot fine-tunability of baselines. This, together with ’vesselFM (from scratch)’ benefiting from our well-evaluated UNet architecture, explains its relatively strong performance.
Appendix E More Detailed Ablation of Data Sources
@@
Figure 11: More slices of exemplary domain randomized images (a), images sampled from our flow matching-based generative model $\mathcal{F}$ (b), and images sampled from Med-DDPM (c). Masks are shown in translucent red.
Appendix F VesselFM’s Utility in (Pre-)Clinical Settings
In (pre-)clinical settings, high-quality annotations for emerging imaging technologies and novel, unstudied structures of interest are often not immediately available. Therefore, clinicians and researchers typically create voxel-level annotations from scratch through labor-intensive manual labeling to train supervised segmentation algorithms, a necessity for automated, large-scale, and accurate analysis. Voxel-level annotations, however, can be far more efficiently obtained via automated pre-segmentation and iterative label refinement. This process is often referred to as bootstrapping. With specialist models failing to bridge domain gaps, vesselFM’s exceptional zero-shot generalization and fine-tunability render it ideal for such applications.
Table 7: Comparison to specialist models trained on individual blood vessel segmentation datasets.
Task Model OCTA [54, 17] BvEM [51] SMILE-UHURA [9] MSD8 [2]
[…]
zero-shot specialist model - TopCoW [59] 39.14 24.03 55.09 52.42 38.71 34.35 4.18 3.45
specialist model - VesSAP [48] 33.31 60.69 50.89 32.75 16.27 25.63 9.40 15.19
specialist model - CSD [10, 6] 21.92 42.98 0.04 0.01 59.17 53.07 0.44 -
vesselFM (ours) 46.94 67.07 67.49 62.04 74.66 75.27 29.69 36.14
one-shot specialist model - TopCoW [59] 64.80 74.39 71.53 78.65 36.52 35.66 17.61 -
specialist model - VesSAP [48] 69.26 76.67 72.31 75.94 27.40 30.50 28.94 35.40
specialist model - CSD [10, 6] 64.62 77.56 68.64 78.96 75.96 78.20 16.09 -
vesselFM (ours) 72.10 83.73 78.27 79.91 76.43 78.36 36.88 48.65
few-shot specialist model - TopCoW [59] 70.28 76.79 65.89 77.76 52.97 47.54 26.25 37.16
specialist model - VesSAP [48] 70.27 77.14 69.36 78.40 46.83 44.31 36.27 44.74
specialist model - CSD [10, 6] 71.11 79.36 59.94 78.61 78.20 78.59 34.28 47.57
vesselFM (ours) 75.70 84.03 78.11 84.54 78.77 79.37 45.04 57.25
Appendix G More Samples From $\mathcal{D}_{\text{drand}}$ and $\mathcal{D}_{\text{flow}}$
We present additional samples from $\mathcal{D}_{\text{drand}}$ and $\mathcal{D}_{\text{flow}}$ in Fig. 11a and b, respectively. In conclusion, one can state that our domain randomized generative pipeline produces a wide variety of image-mask pairs with highly diverse fore- and background geometries and textures, while images sampled from our flow matching-based generative model $\mathcal{F}$ exhibit intensity patterns closely mimicking those of real images in $\mathcal{D}_{\text{real}}$. Please note that Fig. 11c depicts images sampled from the Med-DDPM baseline for comparison.
Appendix H More Masks $\mathcal{M}_{\text{syn}}$
Appendix I Additional Qualitative Zero-Shot Results